# DASPack: Controlled Data Compression for Distributed Acoustic SensingDASPack: Controlled Data Compression for Distributed Acoustic Sensing

### v1

@@ 5.3 Analysis of DASPack’s lossless pipeline
Linear predictive coding plays a complementary role by flattening the residual spectrum. Figures 3 and 4 show that, after LPC, the residuals better approximate white noise, a condition under which arithmetic coding approaches the Shannon limit. The order of the predictor is kept small (typically two or three in time and one in space) to bound the side information that must be stored with each tile. Larger numbers of coefficients were not found to improve compression efficiency and, in fact, tend to degrade it due to the increased volume of side information. Since these coefficients are stored independently for each tile, smaller tile sizes lead to more frequent coefficient storage, which increases overhead and hinders overall compression. However, excessively large tiles may span signals with varying statistical properties, reducing the effectiveness of the linear prediction. A balance between these effects was found by using square tiles of 2000 samples per side, but the optimal choice depends on the data.
Beyond raw performance, DASPack addresses practicalities that often hinder adoption of research prototypes. The core is written in Rust, a language that combines speed with memory-safety guarantees. The native HDF5 backend integrates seamlessly with existing seismic workflows. Python bindings expose the compressor directly to popular data-analysis environments, so that researchers can read and write compressed files with little effort.
@@ 5.5 Applications of DASPack
The fixed-accuracy compression approach allows practitioners to choose a quantisation step that is safely below the minimum signal amplitude of interest. Figs. 8, 9, 10 and 11 demonstrate this principle: near-perfect signal fidelity is achieved at moderate quantisation steps, achieving up to 6×6$\times$6 × compression factors.
@@ 6 Conclusions
Across a diverse set of 15 real-world DAS datasets, DASPack consistently outperformed both general-purpose and domain-specific compression algorithms in terms of compression efficiency and processing speed. File size reductions of up to 3×3$\times$3 × were achieved in lossless mode, with up to 6×6$\times$6 × compression realised with minimal information loss, and 10×10$\times$10 × compression attained with acceptable degradation. It demonstrated real-time capability even under demanding conditions, such as high channel counts and high sampling rates. The tool supports a broad range of acquisition scenarios and integrates easily into existing workflows via HDF5 support and Python bindings.
@@ Acknowledgements.
This work was supported in part by the ”Severo Ochoa Centre of Excellence” program (CEX2019-000928-S-21-2); the European Union NextGenerationEU/PRTR Program under projects PSI (PLEC2021-007875), TREMORS (CPP2021-008869) and the Spanish Ministry of Science, Innovation and Universities under DeeLight project (PID2020-117142GB-I00). Canalink, IslaLink, Telxius, ENAGAS, Teide and Roque de los Muchachos Astronomical Observatories facilitated fibre-optic cable access. Aragon Photonics provided the HDAS interrogator used in some of the experiments.
@@
DASPack is openly available at https://github.com/asleix/daspack, including examples to use the software. We used an existing entropy coding implementation named Constriction, https://github.com/bamler-lab/constriction. For published datasets listed in Table 1, details on open data access can be found in the correponding publications. Unpublished datasets are available from the respective authors upon reasonable request.

### v2

@@ 5.4 Analysis of DASPack’s lossless pipeline
Linear predictive coding plays a complementary role by flattening the residual spectrum. Figures 3 and 4 show that, after LPC, the residuals better approximate white noise, a condition under which arithmetic coding approaches the Shannon limit. Prior to this, subtracting the temporal mean (stored as side information) was found to improve stationarity and thus yield a more effective model. In contrast, removing the spatial mean did not provide any noticeable gain in compression. The order of the predictor is kept small (typically two or three in time and one in space) to bound the side information that must be stored with each tile. Larger numbers of coefficients were not found to improve compression efficiency and, in fact, tend to degrade it due to the increased volume of side information. Since these coefficients are stored independently for each tile, smaller tile sizes lead to more frequent coefficient storage, which increases overhead and hinders overall compression. However, excessively large tiles may span signals with varying statistical properties, reducing the effectiveness of the linear prediction. A balance between these effects was found by using square tiles of 1000 or 2000 samples per side, but the optimal choice depends on the data.
Beyond raw performance, DASPack addresses practicalities that often hinder adoption of research prototypes. The core is written in Rust, a language that combines speed with memory-safety guarantees. The code interface is designed to integrate seamlessly with existing seismic workflows and can be combined with widely used formats such as HDF5 and Zarr. Python bindings expose the compressor directly to popular data-analysis environments, so that researchers can read and write compressed files with little effort. Default parameters were established for the processing steps, accommodating the vast majority of cases and resulting in only marginal variations in compression efficiency, conveniently relieving users from the need to adjust them.
@@ 5.5 Combination with sampling methods
5.6 Comparison and potential combination with deep learning methods
Recent studies have proposed deep learning for DAS data reduction, primarily through subsampling with reconstruction (Ni et al., 2024) and transformer-based autoencoders (Chen et al., 2025). While effective in extracting higher-order structure, these approaches fundamentally differ from DASPack. Ni et al. (2024) approximate missing traces from downsampled data, whereas Chen et al. (2025) encode DAS tiles into latent features, storing decoder weights for signal regeneration. Both strategies produce approximations without explicit error bounds, and may obscure low-amplitude or transient events. In contrast, DASPack provides deterministic guarantees: either exact lossless or fixed-accuracy lossy compression with strictly bounded error.
The computational trade-offs are also substantial. Chen et al. (2025) report $\sim 30$ minutes of GPU training for a one-minute DAS segment, with transfer learning partially mitigating the cost. Ni et al. (2024) achieved 51 seconds of A100 GPU compute to reconstruct a 10-minute record, and $\sim 80$ reconstructions per second on a Raspberry Pi for 1,000 channels. This is equivalent to a $\sim 3\times$ real-time throughput at 25 Hz sampling, with poor scaling for larger deployments. In comparison, DASPack compresses and decompresses in real time on standard CPUs, achieving sustained throughput up to 200 MB $s^{-1}$ using one thread, without requiring GPUs or model storage. This is equivalent to real-time sampling throughput of about 50,000 Hz for the same example of 1,000 channels.
Importantly, downsampling or approximation should not be confused with entropy-based compression. Subsampling discards samples and interpolates missing information, while DASPack re-encodes the full signal bandwidth with user-defined error control. This distinction implies complementarity: DASPack can be applied prior to or after learned reconstruction, providing reliable entropy reduction while leaving room for interpolation-based inpainting. Likewise, DASPack’s controlled quantisation could act as a regulariser for neural models, suppressing noise below application thresholds and potentially accelerating training convergence.
@@ 5.7 Applications of DASPack
The fixed-accuracy compression approach allows practitioners to choose a quantisation step that is safely below the minimum signal amplitude of interest. Figs. 8, 9, 10 and 11 demonstrate this principle: near-perfect signal fidelity is achieved at moderate quantisation steps, achieving up to $6\times$ compression factors.
@@ 6 Conclusions
Across a diverse set of 15 real-world DAS datasets, DASPack consistently outperformed both general-purpose and domain-specific compression algorithms in terms of compression efficiency and processing speed. File size reductions of up to $3\times$ were achieved in lossless mode, with up to $6\times$ compression realised with minimal information loss, and $10\times$ compression attained with acceptable degradation. It demonstrated real-time capability even under demanding conditions, such as high channel counts and high sampling rates. The tool supports a broad range of acquisition scenarios and integrates easily into existing workflows via HDF5 support and Python bindings.
@@ Acknowledgements.
We thank Professor Francesco Grigoli and an anonymous reviewer for the valuable feedback that helped improve the manuscript. The first author thanks Professor Cesar R. Ranero for his support during the realisation of the work. This work was supported in part by the ”Severo Ochoa Centre of Excellence” program (CEX2019-000928-S-21-2); the European Union NextGenerationEU/PRTR Program under projects PSI (PLEC2021-007875), TREMORS (CPP2021-008869), the CSIC contract SAFE with Telxius Cable España, S. L. (ref. 20233069) and the Spanish Ministry of Science, Innovation and Universities under DeeLight project (PID2020-117142GB-I00). Canalink, IslaLink, Telxius, ENAGAS, Teide and Roque de los Muchachos Astronomical Observatories facilitated fibre-optic cable access. Aragon Photonics provided the HDAS interrogator used in some of the experiments.
@@
DASPack is openly available at https://github.com/asleix/daspack, including examples to use the software. We used an existing entropy coding implementation named Constriction, https://github.com/bamler-lab/constriction. For published datasets listed in Table 1, details on open data access can be found in the correponding publications. Unpublished datasets are available from the respective authors upon reasonable request. Data segments used in Figs. 811 are available at https://doi.org/10.20350/digitalCSIC/17586 (Ugalde & Fitchner, 2025).